OpenStudy (anonymous):

Etta deposited 20,000 in a three year certificate of deposit that pays simple interest at a fixed annual rate of 4.9%. What total interest will Etta have earned at the end of three years?

OpenStudy (anonymous):

P = 20,000 r = 4.9% t = 3yrs \[SI = \frac{ P*r*t }{ 100 }\]

OpenStudy (anonymous):

$29.40

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

or do I have to times it by the 3 years

OpenStudy (anonymous):

2940

OpenStudy (anonymous):

ok so just dont divide it by 100?

OpenStudy (anonymous):

\[SI = \frac{ 20,000*4.9*3 }{ 100 } = 200 *4.9*3 = 2940\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Anytime :)